OpenStudy (kj4uts):

A sample of 81 was taken from a population with a standard deviation of 45 kilograms. If the sample mean was 112 kilograms, the mean of any other sample of 81 will be within what range 95% of the time? A. 97 kilograms to 127 kilograms B. 92 kilograms to 132 kilograms C. 102 kilograms to 122 kilograms D. 107 kilograms to 117 kilograms

iYuko (iyuko):

@prettygirl876

jimthompson5910 (jim_thompson5910):

L = xbar - z*s/sqrt(n) U = xbar + z*s/sqrt(n) xbar = sample mean s= standard deviation z = critical value

OpenStudy (kj4uts):

what would (n) be?

OpenStudy (kj4uts):

is 81 the critical value?

jimthompson5910 (jim_thompson5910):

n = sample size

jimthompson5910 (jim_thompson5910):

z = 2 is the critical value because 95% of the data is within 2 std dev of the mean

OpenStudy (kj4uts):

so would it be set up like this:? 112-2*45/sqrt81 112+2*45/sqrt81

OpenStudy (kj4uts):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (kj4uts):

The first one looks like 102 and the second 122 which would be C.?

OpenStudy (kj4uts):

thank you :)